Find the radius of a circle that has an area of 25π yd(squared)
Answer by checkley77(12844)

AREA=PI*R^2 FORMULA FOR AREA OF CIRCLE.
PI*25=PI*R^2
25=R^2
R=SQRT25
R=5 YDS. ANS. FOR THE RADIUS OF THE CIRCLE.